OpenStudy (anonymous):

Chem homework, i think the formula is d=m/v please explain answer too: What is the density of benzene in grams per mL if 5.5L of benzene weighs 4.8kg? (i got 1.14583

OpenStudy (anonymous):

D = m / V = 4.8g / 5.5 ml = 0,872g/ml

OpenStudy (anonymous):

you forgot to convert the units

OpenStudy (anonymous):

1L = 1000 ml, 1kg = 1000g, that cancels them putt, then follow the formula and divide mass/volume :)

OpenStudy (anonymous):

*them out

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

oh no i divided volume to mass instead of mass to volume, thats what it was.

OpenStudy (anonymous):

try it and let me know if i have it right? it should be :) no problem! my pleasure

OpenStudy (anonymous):

yes i redid it and i got the same thing as you. First time posting on this website and i am glad i could find some help so quickly. thanks a lot